# Not All LLM-Generated Data Are Equal: Rethinking data weighting in Text Classification

### v1

@@ Abstract
Synthetic data augmentation via large language models (LLMs) allows researchers to leverage additional training data, thus enhancing the performance of downstream tasks, especially when real-world data is scarce. However, the generated data can deviate from the real-world data, and this misalignment can bring deficient outcomes while applying the trained model to applications. Therefore, we proposed efficient weighted-loss approaches to align synthetic data with real-world distribution by emphasizing high-quality and diversified data generated by LLMs with using merely a little real-world data. We empirically assessed the effectiveness of our method on multiple text classification tasks, and the results showed leveraging our approaches on a BERT-level model robustly outperformed standard cross-entropy and other data weighting approaches, providing potential solutions to effectively leveraging synthetic data from any suitable data generator for model training.
@@ 1 Introduction
The quantity and quality of data play a significant role in many tasks of Natural Language Processing (NLP). However, due to the scarcity of data in a particular domain for a specific task, we may need expertise to collect such data, resulting in budget limitations. Fortunately, Large language models (LLMs) provide a practical solution to this problem. LLMs, such as GPT series (Brown et al., 2020; OpenAI, 2022; OpenAI et al., 2024), can be leveraged to generate synthetic data that mimics real-world examples, thereby enriching the training set (Wang et al., 2023). (Taori et al., 2023), and other works have shown the capability of using LLM-generated data for the downstream tasks, and it seems to be a new cut-in solution to any NLP downstream tasks. However, training models with LLM-generated data can lead to drawbacks such as model collapse (Shumailov et al., 2023; Dohmatob et al., 2024), tail phenomena, reinforcing LM biases (Wang et al., 2023). Moreover, based on our empirical study, the performance of models trained on synthetic data without proper processing can be lower than models trained on much smaller real-world data (Sec. 3.1), highlighting the uncertainty of using LLM-generated data.
Previous works took data filtering strategy to get high quality or variant data (Dubey et al., 2024; MetaAI, 2024; Chiang et al., 2023; West et al., 2022). Still, this strategy was mostly human-crafted or needed efforts to train a judge model, and most importantly, filtering strategies abandoned the potential of the filtered data that may contribute to the final performance. In contrast, data weighting approaches leverage all the training data, including augmented and biased data, but prioritize data by giving nonuniform weights to the loss of each data point. For example, Focal-Loss (Lin et al., 2017) prioritized more diverse data, and Hu et al. (2019) optimized the weights of training samples so that the model performs best on a small real-world dataset. It is worth noting that while Hu et al. (2019) steered the training toward higher performance, using this method in a large-scale scenario seems infeasible because it took the weights as learnable parameters, the number of which increased as the training set scale. Thus, inspired by the objective of (Hu et al., 2019), we introduce two novel, efficient, automatic weighted-loss approaches: Importance Loss (IMP-Loss) and Dynamic Importance Loss (DIMP-Loss), which are designed to closely align the distribution of synthetic data with that of real-world data. Furthermore, both IMP-Loss and DIMP-Loss incorporate mechanisms as quality-checkers and diversity-checkers, assigning higher quality and more diverse data points with greater weight. In other words, these methods prioritize data points that are more relevant and more informative to target downstream tasks, thereby reducing the impact of less valuable data on the fine-tuning model.
@@ 2.1 Cross Entropy Loss on Real-World Dataset
In supervised learning for text classification tasks, we consider a real-world dataset $D_{P}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{M}$ comprising $M$ samples. Each pair $(\mathbf{x}_{i},y_{i})$ is drawn independently and identically distributed (i.i.d.) from the joint distribution $P(\mathcal{X},\mathcal{Y})$, where $\mathbf{x}_{i}$ represents the input sample and $y_{i}\in\mathcal{Y}=\{1,2,\ldots,C\}$ is the corresponding class label.
This setup forms the basis for training models using the empirical cross-entropy (CE-Loss) as the loss, a standard objective function in such tasks. The CE-Loss over the entire dataset $D_{P}$ is calculated as follows:
@@ 2.2 Weighted Cross Entropy Loss (WCE-Loss)
Here, $w_{i}$ represents the weight assigned to the $i$-th data point $(\mathbf{x}_{i},y_{i})$. A higher weight $w_{i}$ assigned to a data point $(\mathbf{x}_{i},y_{i})$ indicates that the data point is considered more important for the training process, thereby having a greater influence on the model’s learning or adjustment of parameters.
@@
Additionally, a series of meta-learning approaches addressed these challenges by leveraging bi-level optimization to dynamically adjust weights based on each instance’s contribution to the meta-set from the real world. These methods handled class imbalance, noisy labels, and augmented data by reweighting these instances based on their gradient direction or model outputs, providing a flexible mechanism for weighting data points (Ren et al., 2018; Hu et al., 2019). While effective, meta-learning-based approaches were time-consuming and computationally expensive, making them difficult to scale up to larger datasets or complex models. In contrast, our methods share the same objective of optimizing performance on real-world data but achieve it without meta-learning. This makes it more computationally efficient and scalable while still maintaining high performance.
@@ 3.1 Uncertainty of LLM-Generated Data Performance
Our empirical study, shown in Table 1, demonstrates notable variability in the performance of CE-Loss on LLM-generated datasets. Specifically, on the Financial (Malo et al., 2014) and MRPC (Wang et al., 2018) benchmarks, CE-Loss on large LLM-generated datasets ($>$ 3k samples) performs worse than training on small real-world datasets, which contain only around 200-400 samples. Conversely, for the Twitter Irony (Van Hee et al., 2018) benchmark, CE-Loss on LLM-generated data improves accuracy. This variability underscores the uncertainty associated with using CE-Loss on LLM-generated data. These findings are consistent with another research (West et al., 2022), showing that when using CE-Loss, without proper filtering, LLM-generated data may lead to decent results on downstream tasks, even though its size is considerably larger than that of real-world data.
@@ 3.2 Potential of LLM-Generated Data: Model-Based Information Measurement
We employ information-theoretic metrics to evaluate the uncertainty within the conditional distributions of the real-world data and LLM-generated data. A higher conditional entropy indicates more significant uncertainty given an input $\mathbf{x}$, suggesting various outcomes. We estimate this by fine-tuning a BERT model (Devlin et al., 2019) on both datasets separately. Higher conditional entropy is often associated with greater diversity within the dataset, reflecting a broader range of information that the model must learn to predict accurately. The conditional KL divergence111it is also called conditional divergence or conditional relative entropy quantifies the difference between two conditional distributions, $P(y|\mathbf{x})$ and $Q(y|\mathbf{x})$, showing how well a model trained on one dataset describes another.
We show these metrics for a Financial benchmark scenario. The real-world dataset $D_{P}$ exhibits significantly lower conditional entropy ($\mathbb{H}(P)=0.0365$) compared with the LLM-generated dataset $Q$ ($\mathbb{H}(Q)=0.2299$), indicating that $D_{Q}$ is more diverse. Furthermore, the conditional KL divergence from $P$ to $Q$ ($D_{\mathrm{KL}}(Q||P)=1.8781$) is much greater than it from $Q$ to $P$ ($D_{\mathrm{KL}}(P||Q)=0.444$), suggesting that models trained on real-world data struggle to capture the complexity of the synthetic dataset. Those models trained on the synthetic dataset are relatively efficient, requiring fewer additional nits on average to encode samples from the real-world distribution $P$. This difference, along with the results from Sec. 3.1 highlights that, although the synthetic dataset contains some points that are less representative of the real-world distribution, it still includes a substantial proportion of relevant data points. This analysis indicates potential for improving modeling techniques to utilize the rich, informative content of LLM-generated data.
@@ 3.3 Problem Formulation
In this study, we devise a weighted loss function that transforms CE-Loss from an LLM-generated data distribution $Q$ to match the real-world data distribution $P$. We assumed the dataset $D_{Q}$ is i.i.d. and the LLM can approximate the real-world input distribution $P(\mathbf{x})$ through strategic prompting, effectively simulating $Q(\mathbf{x})\approx P(\mathbf{x})$. For instance, by using system prompts like ”Now you are a journalist writing news articles,” it can produce synthetic texts that closely mimic authentic news articles. Lastly, We use a small set $D_{P^{\prime}}$, approximately 200-400 samples from real-world datasets, to facilitate the alignment process. These samples are i.i.d. from the distribution $P$. We use $P^{\prime}$ as the probability function representing this small real-world dataset.
@@ 4.1 IMP-Loss: Transformation from $Q$ to $P$
In order to achieve convergence to the real-world data distribution $P$, we applied WCE-loss. Inspired by the Monte Carlo method of Importance Sampling (Hesterberg, 1995), used to estimate expectation values from a source distribution to a target distribution, we design the weight function as follows:
@@
The approximation in the penultimate step is based on the assumption stated in Sec. 3.3: the LLM can simulate the real-world input distribution through careful and appropriate prompting. This transformation ensures that the WCE-Loss function effectively aligns the synthetic distribution $Q$ with the real-world distribution $P$.
@@ 4.2 DIMP-Loss: Which data point causes the model to be closest to $P$?
where $\theta_{t}$ represents the model parameters at optimisation step $t$. Consider a one-step optimization algorithm $f$ (e.g. SGD), where $\theta_{t+1}\leftarrow f(\theta_{t},\{(\mathbf{x}^{\prime},y^{\prime})\})$. The algorithm updates the model parameters $\theta_{t}$ using $(\mathbf{x}^{\prime},y^{\prime})$ to obtain the new parameters $\theta_{t+1}$ after one optimization step. The Eq. 7 means the data point $(\mathbf{x}^{*},y^{*})$ is the optimal data point in $D_{Q}$ that leads to the lowest conditional cross-entropy after one update step. Specifically, it identifies which data point is used for training results in the model parameters that yield the model closest to the real-world distribution $P$.
@@
Notably, by applying Bayes’ rule, we derive the second line of Eq. 8 (see Appendix C for details), showing a more feasible calculation approach. This requires evaluating only two models for each data point in $D_{Q}$: the denominator $\hat{P}(y|\mathbf{x};\theta_{t})$ is the current model in step $t$, and the numerator $\hat{P}(y|\mathbf{x};\theta_{t},D_{P^{\prime}})$ would require additional training on $D_{P^{\prime}}$. To simplify, we approximate $\hat{P}(y|\mathbf{x};\theta_{t},D_{P^{\prime}})$ with $\hat{P^{\prime}}(y|\mathbf{x})$, the probability estimated from $D_{P^{\prime}}$ as in (Deng et al., 2023).
The approximation of Eq. 8 is then utilized as the weight in our loss function. Consequently, if a data point brings the model closer to the real-world distribution $P$, its corresponding weight will be higher, thus having a greater impact on the model’s training. Thus, we define the Dynamic Importance Loss (DIMP-Loss) $\mathcal{L}_{\text{DIMP}}(\theta_{t},D_{Q})$ as:
@@
DIMP-Loss can be interpreted as an upper bound on the regularized empirical distilled risk (Menon et al., 2021; Wang et al., 2022), where the ”teacher” model is the quality checker. The regularizer is a maximum entropy term (Pereyra et al., 2017) designed to prevent overconfidence in output distribution. In this context, DIMP-Loss can also be viewed as a form of knowledge distillation, where the knowledge from a model is trained on a small amount of real-world data. The objective is to align the predicted distribution $\hat{P}_{\theta}$ with $P^{\prime}$ while promoting higher entropy in $\hat{P}_{\theta}$ to avoid overly confident predictions.
@@ 4.3 Quality and Diversity Checkers in IMP-Loss and DIMP-Loss
According to both Eq. 6 and 9, a higher weight value in either loss means that the data point will have a greater influence on the model.
the Quality Checker ($\hat{P^{\prime}}(y_{i}|\mathbf{x}_{i})$) assesses the likelihood of a data point sampled from the real-world distribution $P$. Higher values indicate the data point is, highly relevant, and unambiguous for the real-world distribution $P$.
The Diversity Checker differs in the two losses, $\hat{Q}(y_{i}|\mathbf{x}_{i})$ for IMP-Loss, and $\hat{P}(y_{i}|\mathbf{x}_{i};\theta_{t})$ for DIMP-Loss. In the context of IMP-Loss, a low Diversity Checker value ($\hat{Q}(y_{i}|\mathbf{x}_{i})$) suggests the data point contains a high amount of information within the LLM-generated dataset $D_{Q}$, because a redundant data point in $D_{Q}$ will have a high probability, indicating less diversity. Hence, the Diversity Checker serves as an indicator of diversity from the perspective of the LLM-generated distribution. In contrast, for DIMP-Loss, a low Diversity Checker value ($\hat{P}(y_{i}|\mathbf{x}_{i};\theta_{t})$) implies the data point is challenging to be learnt in previous steps, departing from the data points the model has already learnt. Thus, Diversity Checker of DIMP-Loss reflects diversity from the perspective of a model. This distinction highlights how each loss function prioritizes different aspects of data diversity during training. We simulated defect and redundant situations for further exploration in the Appendix. G.3.

### v2

@@ Abstract
Synthetic data augmentation via Large Language Models (LLMs) allows researchers to leverage additional training data, thus enhancing the performance of downstream tasks, especially when real-world data is scarce. However, the generated data can deviate from the real-world data, and this misalignment can bring about deficient results while applying the trained model to applications. Therefore, we proposed efficient weighted-loss approaches to align synthetic data with real-world distribution by emphasizing high-quality and diversified data generated by LLMs using merely a tiny amount of real-world data. We empirically assessed the effectiveness of our methods on multiple text classification tasks, and the results showed that leveraging our approaches on a BERT-level model robustly outperformed standard cross-entropy and other data weighting approaches, providing potential solutions to effectively leveraging synthetic data from any suitable data generator.
@@ 1 Introduction
The quantity and quality of data play a significant role in many tasks of Natural Language Processing (NLP). However, due to the scarcity of data in a particular domain for a specific task, we may need expertise to collect such data, resulting in budget limitations. Fortunately, Large Language Models (LLMs) provide a practical solution to this problem. LLMs, such as GPT series (Brown et al., 2020; OpenAI, 2022; OpenAI et al., 2024), can be leveraged to generate synthetic data that mimics real-world examples, thereby enriching the training set (Wang et al., 2023). Taori et al. (2023), and other works (Ye et al., 2022; West et al., 2022; Li et al., 2023) have shown the capability of using LLM-generated data for the downstream tasks, and it seems to be a new cut-in solution to any NLP downstream tasks. However, training models with LLM-generated data can lead to drawbacks such as model collapse (Shumailov et al., 2023; Dohmatob et al., 2024), tail phenomena, reinforcing LM biases (Wang et al., 2023). Moreover, based on our empirical study, the performance of models trained on synthetic data without proper processing can be lower than models trained on much smaller real-world data (Sec. 3.1), highlighting the uncertainty of using LLM-generated data.
Previous works took filtering strategy to get high quality or variant data (Dubey et al., 2024; MetaAI, 2024; Chiang et al., 2023; West et al., 2022; Meng et al., 2022; 2023). Still, this strategy was mostly human-crafted or needed efforts to train a judge model, and most importantly, filtering strategies abandoned the potential of the filtered data that may contribute to the final performance. In contrast, data weighting approaches leverage all the training data, including augmented and biased data, but prioritize data by giving nonuniform weights to the loss of each data point. For example, Focal-Loss (Lin et al., 2017) prioritized more diverse data; Hu et al. (2019) and SunGen (Gao et al., 2023) optimized the weights of training samples so that the model performs best on a small real-world dataset. It is worth noting that while Hu et al. (2019) and SunGen steered the training toward higher performance, using these methods in a large-scale scenario seems infeasible because the weights are regarded as learnable parameters, the number of which increases as the training set grows.
Thus, inspired by the objective of Hu et al. (2019), we introduce two novel, efficient, automatic weighted-loss approaches: Importance Loss (IMP-Loss) and Dynamic Importance Loss (DIMP-Loss), which are designed to closely align the distribution of synthetic data with that of real-world data. Furthermore, both IMP-Loss and DIMP-Loss incorporate mechanisms as quality-checkers and diversity-checkers, assigning higher quality and more diverse data points with greater weight. In other words, these methods prioritize data points that are more relevant and more informative to target downstream tasks, thereby reducing the impact of less valuable data on the fine-tuning model.
@@ 2.1 Cross Entropy Loss on Real-World Dataset
In supervised learning for text classification tasks, we consider a real-world dataset $D_{P}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{M}$ comprising $M$ samples. Each pair $(\mathbf{x}_{i},y_{i})$ is drawn independently and identically distributed (i.i.d.) from the joint distribution $P(\mathcal{X},\mathcal{Y})$, where $\mathbf{x}_{i}$ represents the input sample and $y_{i}\in\mathcal{Y}=\{1,2,\ldots,C\}$ is the corresponding class label. This setup forms the basis for training models using the empirical cross-entropy loss (CE-Loss), a standard loss function in such tasks. The CE-Loss over the entire dataset $D_{P}$ is calculated as follows:
@@ 2.2 Weighted Cross Entropy Loss (WCE-Loss)
Here, $w_{i}$ represents the weight assigned to the $i$-th data point $(\mathbf{x}_{i},y_{i})$. A higher weight $w_{i}$ assigned to a data point $(\mathbf{x}_{i},y_{i})$ indicates the data point has a greater influence on the model’s learning or adjustment of parameters, thereby being considered more important for the training process.
@@
Additionally, a series of meta-learning approaches addressed these challenges by leveraging bi-level optimization to dynamically adjust weights based on each instance’s contribution to the meta-set from the real world. These methods handled class imbalance, noisy labels, and augmented data by reweighting these instances based on their gradient direction or model outputs, providing a flexible mechanism for weighting data points (Ren et al., 2018; Hu et al., 2019; Gao et al., 2023). While effective, meta-learning-based approaches were computationally expensive, making them difficult to scale up to larger datasets or complex models. In contrast, our methods share the same objective of optimizing performance on real-world data but achieve it without meta-learning. This makes it more computationally efficient and scalable while still maintaining high performance.
@@ 3.1 Uncertainty of LLM-Generated Data Performance
Our empirical study, shown in Table 1, demonstrates notable variability in the performance of CE-Loss on LLM-generated datasets. Specifically, on the Financial (Malo et al., 2014) and MRPC (Wang et al., 2018) benchmarks, CE-Loss on large LLM-generated datasets ($>$ 3k samples) performs worse than training on small real-world datasets, which contain only around 200-400 samples. In contrast, CE-Loss in LLM-generated data improves accuracy for the Twitter Irony (Van Hee et al., 2018) benchmark. This variability underscores the uncertainty associated with using CE-Loss on LLM-generated data. These findings are consistent with another research (West et al., 2022), showing that when using CE-Loss, without proper filtering, LLM-generated data may lead to decent results on downstream tasks, even though its size is considerably larger than that of real-world data.
@@ 3.2 Potential of LLM-Generated Data: Model-Based Information Measurement
We employ information-theoretic metrics to evaluate the uncertainty within the conditional distributions of real-world data and LLM-generated data. A higher conditional entropy indicates a more significant uncertainty given an input $\mathbf{x}$, suggesting various outcomes. We estimate this by separately fine-tuning a BERT model (Devlin et al., 2019) on both datasets. Higher conditional entropy is often associated with greater diversity within the dataset, reflecting a broader range of information that the model must learn to predict accurately. The conditional KL divergence [^1] It is also called conditional divergence or conditional relative entropy. quantifies the difference between two conditional distributions, $P(y|\mathbf{x})$ and $Q(y|\mathbf{x})$, showing how well a model trained on one dataset describes another.
We show these metrics for a financial benchmark scenario. The real-world dataset $D_{P}$ exhibits significantly lower conditional entropy ($\mathbb{H}_{P}(y|\mathbf{x})=0.0365$) compared with the LLM-generated dataset $Q$ ($\mathbb{H}_{Q}(y|\mathbf{x})=0.2299$), indicating that $D_{Q}$ is more diverse. Furthermore, the conditional KL divergence from $P$ to $Q$ ($D_{\mathrm{KL}}(Q||P)=1.8781$) is much greater than it from $Q$ to $P$ ($D_{\mathrm{KL}}(P||Q)=0.444$), suggesting that models trained on real-world data struggle to capture the complexity of the synthetic dataset. Those models trained on the synthetic dataset are relatively efficient, requiring fewer additional nits on average to encode samples from the real-world distribution $P$. This difference, along with the results from Sec. 3.1 highlights that, although the synthetic dataset contains some points that are less representative of the real-world distribution, it still includes a substantial proportion of relevant data points. This analysis indicates the potential to improve modeling techniques to utilize LLM-generated data’s rich, informative content.
@@ 3.3 Problem Formulation
In this study, we devise a weighted loss function that transforms CE-Loss from an LLM-generated data distribution $Q$ to match the real-world data distribution $P$. We assumed the dataset $D_{Q}$ is i.i.d. and the LLM can approximate the real-world input distribution $P(\mathbf{x})$ through strategic prompting, effectively simulating $Q(\mathbf{x})\approx P(\mathbf{x})$. For example, by using system prompts like ”Now you are a journalist writing news articles,” it can produce synthetic texts that closely mimic authentic news articles. Lastly, we use a small set $D_{P^{\prime}}$, approximately 200-400 samples from real-world datasets, to facilitate the alignment process. These samples are i.i.d. from the distribution $P$. We use $P^{\prime}$ as the probability function representing this small real-world dataset.
@@ 4.1 IMP-Loss: Transformation from $Q$ to $P$
To achieve convergence to the real-world data distribution $P$, we applied WCE-loss. Inspired by the Monte Carlo method of Importance Sampling (Hesterberg, 1995), used to estimate expectation values from a source distribution to a target distribution, we design the weight function as follows:
@@
The approximation in the penultimate step is based on the assumption stated in Sec. 3.3: the LLM can simulate the real-world input distribution through careful and appropriate prompting. This transformation ensures that the WCE-Loss effectively aligns $Q$ with $P$.
@@ 4.2 DIMP-Loss: Which data point causes the model to be closest to $P$?
where $\theta_{t}$ represents the model parameters at optimization step $t$. Consider a one-step optimization algorithm $f$ (e.g. SGD), where $\theta_{t+1}\leftarrow f(\theta_{t},\{(\mathbf{x}^{\prime},y^{\prime})\})$. The algorithm updates the model parameters $\theta_{t}$ using $(\mathbf{x}^{\prime},y^{\prime})$ to obtain the new parameters $\theta_{t+1}$ after one optimization step. The Eq. 7 means the data point $(\mathbf{x}^{*},y^{*})$ is the optimal data point in $D_{Q}$ that leads to the lowest conditional cross-entropy after one update step. Specifically, it identifies which data point is used for training results in the model parameters that yield the model closest to the real-world distribution $P$.
@@
Notably, by applying Bayes’ rule, we derive the right-hand side of Eq. 8 (see Appendix C for details), showing a more feasible calculation approach. This requires evaluating only two models for each data point in $D_{Q}$: the denominator $\hat{P}(y|\mathbf{x};\theta_{t})$ is the current model in step $t$, and the numerator $\hat{P}(y|\mathbf{x};\theta_{t},D_{P^{\prime}})$ would require additional training on $D_{P^{\prime}}$. To simplify, we approximate $\hat{P}(y|\mathbf{x};\theta_{t},D_{P^{\prime}})$ with $\hat{P^{\prime}}(y|\mathbf{x})$, the probability estimated from $D_{P^{\prime}}$ as in Deng et al. (2023).
The approximation of Eq. 8 is then utilized as the weight in our loss function. Consequently, if a data point brings the model closer to the real-world distribution $P$, its corresponding weight will be higher, thus having a greater impact on training. Thus, we define the Dynamic Importance Loss (DIMP-Loss) $\mathcal{L}_{\text{DIMP}}(\theta_{t},D_{Q})$ as:
@@
DIMP-Loss can be interpreted as an upper bound on the regularized empirical distilled risk (Menon et al., 2021; Wang et al., 2022), where the ”teacher” model is the quality checker. The regularizer is a maximum entropy term designed to prevent overconfidence in output distribution (Pereyra et al., 2017). In this context, DIMP-Loss can also be viewed as a form of knowledge distillation, where the knowledge from a model is trained on a small amount of real-world data. The objective is to align the predicted distribution $\hat{P}_{\theta}$ with $P^{\prime}$ while promoting higher entropy in $\hat{P}_{\theta}$ to avoid overly confident predictions.
@@ 4.3 Quality and Diversity Checkers in IMP-Loss and DIMP-Loss
According to both Eq. 6 and 9, a high weight means the data point significantly influences the model. The Quality Checker ($\hat{P^{\prime}}(y_{i}|\mathbf{x}_{i})$) assesses the likelihood of a data point sampled from the real-world distribution $P$. Higher values indicate the data point is, highly relevant, and unambiguous for the real-world distribution $P$.
The Diversity Checker differs in the two losses, $\hat{Q}(y_{i}|\mathbf{x}_{i})$ for IMP-Loss, and $\hat{P}(y_{i}|\mathbf{x}_{i};\theta_{t})$ for DIMP-Loss. In the context of IMP-Loss, a low Diversity Checker value $\hat{Q}(y_{i}|\mathbf{x}_{i})$ suggests the data point contains a high amount of information within the LLM-generated dataset $D_{Q}$, because a redundant data point in $D_{Q}$ will have a high probability, indicating less diversity. Hence, it serves as an indicator of diversity from the perspective of the LLM-generated distribution. In contrast, for DIMP-Loss, a low Diversity Checker value $\hat{P}(y_{i}|\mathbf{x}_{i};\theta_{t})$ implies the data point is challenging to be learned in previous steps, departing from the data points the model has already learned. Thus, Diversity Checker of DIMP-Loss reflects diversity from the perspective of a model. This distinction highlights how each loss function prioritizes different aspects of data diversity during training. We simulated defect and redundant situations for further exploration in the Appendix. G.3.